OpenStudy (anonymous):

if i have a right triangle but only know the hypotenuse which is 11 how do i find x and y (the legs) ?

OpenStudy (jadeishere):

11^2= x^2 + y^2 121 = x^2 + y^2

OpenStudy (mathmale):

Welcome to OpenStudy, Anny! Key to understanding the solution to your math problem is knowing the Pythagorean Theorem. IF and only if you have a right triangle, the square of the hypotenuse is equal to the SUM of the SQUARES of the other two sides (which we call the "legs": Borrowing from Jade (who is here now), (hypotenuse)^2= x^2 + y^2. All you can do here is to let the hyp = 11, obtaining Jade's 121 = x^2 + y^2. For any x value you plug in, you can then calculate the corresponding y value.

OpenStudy (anonymous):

thank you both (: i understand that part, i dont understand what to do after i get 121 = x^2 + y^2 ? do i just leave it or what happens after?

OpenStudy (anonymous):

A kite costs $9.60. Today is is on sale at one third off. What is the sale price of the kite?